The length of a rectangle is four times its width.
If the perimeter of the rectangle is 60yd , find its area.

Answer by mananth(16949)

The length of a rectangle is four times its width.
If the perimeter of the rectangle is 60yd , find its area
Let the width be x
four times width =4x
60= 2*(l+b)
60= 2*(4x+x)
60= 10x
x=6
Length= 4x= 4*6=24
Area of rectangle = L*W
24*6= 144 sq. yards

M Ananth______ mananth@hotmail.com